# Denoising Diffusion Planner: Learning Complex Paths from Low-Quality Demonstrations

### v1

@@ Abstract
Denoising Diffusion Probabilistic Models (DDPMs) are powerful generative deep learning models that have been very successful at image generation, and, very recently, in path planning and control. In this paper, we investigate how to leverage the generalization and conditional-sampling capabilities of DDPMs to generate complex paths for a robotic end effector. We show that training a DDPM with synthetical and low-quality demonstrations is sufficient for generating nontrivial paths reaching arbitrary targets and avoiding obstacles. Additionally, we investigate different strategies for conditional sampling combining classifier-free and classifier-guided approaches. Eventually, we deploy the DDPM in a receding-horizon control scheme to enhance its planning capabilities. The Denoising Diffusion Planner is experimentally validated through various experiments on a Franka Emika Panda robot. Videos of our experiments can be found here.
@@ 1 Introduction
Path planning is a fundamental aspect of almost any robotic task. It can be defined as finding a collision-free trajectory between a starting state and a goal state, such that the system can be driven to the desired target by tracking this trajectory using a low-level controller. Designing algorithms to find these trajectories is difficult, because robotic environments are typically high-dimensional, complex, and dynamic.
@@
Instead, the problem of path planning can be addressed through the paradigm of planning as inference (PAI) [15, 16]. Differently from explicit graph search or autoregressive prediction, the future is modeled as a joint probability distribution over states, actions and rewards. Sampling from this distribution yields a sequence of states and actions that represent a possible trajectory. Retrieving just any possible trajectory is not very useful by itself, but if this trajectory is known to meet certain conditions, like obstacle avoidance or achieving certain rewards, this sample can be used as a plan. In PAI, such plan can be sampled from a conditional probability distribution of possible futures. This process of conditional sampling is key for generating a plan that exhibits desirable behavior.
@@
Figure 1: Overview of the DDP pipeline from low-quality and synthetic demonstration to to real-world tasks.
The generative capabilities of DDPMs can also be used to generate plans. Janner et al. [20] trained DDPMs on an offline collection of state-action sequences, with associated rewards. In [20] it was shown that DDPMs exhibit good generalization and long-horizon planning abilities, even for sparse rewards. Furthermore, they showed that DDPMs can generate novel trajectories by locally stitching together sequences from the training distribution. Moreover, Ajay et al. [21] found that DDPMs can generate plans that meet multiple conditions even though these conditions were never met simultaneously in the training data. These properties of DDPMs make them potentially powerful and versatile planners. Real-world demonstrations of using DDPMs for robot control have shown promising results [22, 23, 24], which inspires this work to further investigate the novel aspects of using DDPMs to solve robotic planning problems. Related work can be found in [24, 25, 22]. Carvalho et al. [24] propose a motion planner in the joint space using DDPMs. They use expert examples produced by an optimal motion planning algorithm to train a DDPM. During inference, they use the gradients of a cost function to bias the samples towards regions of low cost at every denoising step. Saha et al. [25] expand on this work by using an ensemble of cost functions to create a planner that can generate collision-free trajectories in a variety of environments. Chi et al. [22] use DDPMs to directly learn a policy that takes visual observations as its input and produces an action sequence as an output. In their work, they introduce closed-loop planning with DDPMs by implementing a receding-horizon control scheme. DDPM can also be used for safety-critical planning in combination with control barrier functions [26] to generate optimal and safe plans [27, 28].
@@
    we propose a general approach to designing a path planner for an end effector using a DDPM without relying on a dynamic model of the robot;
@@
    we rely only on low-quality and synthetic demonstrations, i.e., straight paths, to train the DDPM;
  3. 3.
    we validate the proposed path planner through real-world experiments.
@@ 3 Denoising Diffusion Planner
Path planning is defined as finding a collision-free motion from a starting configuration to a goal configuration [32]. In this work, each path is discretized into a sequence of poses assumed to be equidistant in time, although we are not limited to that. We indicate a path with $\bm{\tau}=\left(H_{0},H_{1},...,H_{T}\right)$, where $H$ is an element of the Lie-group SE(3) [33] which encodes the position and orientation of a frame with respect to a chosen reference frame. In particular, a pose $H_{t}$ can be described by a matrix that is composed of a rotation matrix $R_{t}$, encoding the orientation, and a vector $\bm{p}_{t}$, encoding the position [33]:
@@ 4 Experimental Design
Our code can be found here.
@@ 4.2 Real-world Experiments
where $d_{SE(3)}\left(H_{t},H_{g}\right)=\lVert\bm{p}_{t}-\bm{p}_{g}\rVert^{2}_{2}+\lVert LogMap\left(\bm{R}^{T}_{t}\bm{R}_{g}\right)\rVert^{2}_{2}$. The $LogMap(\cdot)$ is a function from Lie Theory [33] that can be used to quantify the difference between two orientations. The collision cost is given by:
@@
with the weighting factors $w_{ee}$ and $w_{c}$. Thirdly, we experiment with a novel way of sampling through combining classifier-free sampling with cost-guided sampling. In total, this results in five combinations for conditional sampling: (i) classifier-free guidance with dense rewards, (ii) classifier-free guidance with sparse rewards, (iii) cost guidance using (16), (iv) classifier-free guidance with dense rewards combined with cost guidance (using only the end-effector cost in (14)), and (v) classifier-free guidance with sparse rewards combined with cost guidance (using only the end-effector cost in (14)).
@@ 5 Results and Discussions
In this section, we show the results achieved with DDP. In particular, we study the impact of the choice of horizon and returns when conditionally-generating paths (Section 5.1), the effect of repeatedly inpainting the goal state (Section 5.2), and the performance of the DDP in a real-world robotics task with different conditional-sampling approaches (Section 5.3).
@@ 5.1 Impact of Horizon and Returns
Since the training dataset contains paths in the entire workspace, even inside obstacles, the avoidance of obstacles is entirely achieved through conditional sampling. When using classifier-free guidance, these conditions are learned from the training data, which only consists out of straight lines. The extent to which the planner can generate paths that avoid an obstacle in between the start pose and goal pose depends on the DDPM’s generalization capabilities, but also on the hyperparameters used in inference. In particular, we focus our studies on two important hyperparameters: (i) the planning horizon and (ii) the (expected) return on which the sample is conditioned. We consider a simple scenario in which the DDP is used in open-loop configuration to inpaint a path between a starting pose and a goal pose, while an obstacle is positioned in between. Figure 5 shows simulations of the paths for various combinations of horizon and return values. For the model trained on sparse rewards, only a return value of zero is considered, since this is the only value that represents a collision-free path. Therefore, we effectively only sweep over the horizon in the sparse-reward case. From Figure 5, it is evident that in all configurations, coherent and nontrivial paths are generated. These paths start at the intended starting pose and tend towards the goal pose, even for horizons that are much longer than seen in training (see Table 1). The paths that were generated using sparse rewards are close to the obstacles (Figure 5(b)), whereas for dense rewards, the distance to the obstacle is strongly influenced by the rewards (Figure 5(a)).
@@ 5.3 Real-world Results
Figure 7 shows some frames of the video recording of the real-world experiments. Videos of our experiments can be found here. We compare five different conditional-sampling strategies (see Section 4.2).
The real-world experiments validate that the DDP, trained only on low-quality and synthetic data, is capable of repeatedly generate collision-free path in both test scenarios with different state and goal poses. The DDP is also extremely flexible framework that can exploit different configurations and conditional sampling strategies.
With reference to Figure 7, the DDP trained with sparse rewards (sparse) resulted in paths much closer to the obstacles compared to the DDP trained dense rewards (dense), regardless of the inclusion of cost guidance (cg). However, in the second scenario with two obstacles, using only classifier-free guidance with dense rewards often results in lingering in place before crossing from one side of the smaller obstacle to the other. It is worth mentioning that our reward functions, i.e., sparse and dense, only penalizes collisions or distance to obstacles, respectively, and reaching the goal state is controlled by the fixing the value of the final state(s) of the generated trajectory (see Section 5.2). Therefore, we are not explicitly asking the DDP, except when usiong the cost guidance, for the shortest collision-free path from state to goal state, but simply for a "good" collision-free path.
The DDP with classifier-free guidance with dense rewards can be improved by using the cost function guidance (dense + cg). In particular, in the scenario with two obstacles, the DDP without cost function guidance would often linger in place to avoid moving closer to the obstacles. Adding the cost guidance allows for solving this shortcoming by pushing the poses to be close to the goal pose.
@@ 6 Conclusion and Future Work
In this work, we proposed DDP, an approach to designing a path planner for a robotic end effector using DDPMs that does not use expert demonstrations, but only low-quality and synthetic data. Using only straight lines, the DDP can produce complex paths that avoid obstacles by utilizing the generalization capabilities of the DDPM and different conditional-sampling strategies. Additionally, deploying the DDP in a closed-loop receding-horizon control scheme and continuously update the plan allowes for the DDP to reach difficult targets that cannot be reached in an open-loop configuration.

### v2

@@ Abstract
Denoising Diffusion Probabilistic Models (DDPMs) are powerful generative deep learning models that have been very successful at image generation, and, very recently, in path planning and control. In this paper, we investigate how to leverage the generalization and conditional sampling capabilities of DDPMs to generate complex paths for a robotic end effector. We show that training a DDPM with synthetic and low-quality demonstrations is sufficient for generating nontrivial paths reaching arbitrary targets and avoiding obstacles. Additionally, we investigate different strategies for conditional sampling combining classifier-free and classifier-guided approaches. Eventually, we deploy the DDPM in a receding-horizon control scheme to enhance its planning capabilities. The Denoising Diffusion Planner is experimentally validated through various experiments on a Franka Emika Panda robot. Videos of our experiments can be found here.
@@ 1 Introduction
Path planning is a fundamental aspect of almost any robotic task. It consists in finding a collision-free path between a starting state and a goal state, such that the system can be driven to the desired target by tracking this path using a low-level controller. Designing path planning algorithms is a difficult problem, because robotic environments are typically high-dimensional, complex, and dynamic.
@@
Instead, the problem of path planning can be addressed through the paradigm of planning as inference (PAI) [15, 16]. Differently from explicit graph search or autoregressive prediction, the future is modeled as a joint probability distribution over states, actions and rewards. Sampling from this distribution yields a sequence of states and actions that represent a possible path. Retrieving just any possible path is not very useful by itself, but if this path is known to meet certain conditions, like obstacle avoidance or achieving certain rewards, this sample can be used as a plan. In PAI, such plan can be sampled from a conditional probability distribution of possible futures. This process of conditional sampling is key to generate a plan that exhibits desirable behavior.
@@
Figure 1: Overview of the DDP pipeline from low-quality and synthetic demonstration to real-world tasks.
The generative capabilities of DDPMs can also be used to generate plans. Janner et al. [20] trained DDPMs on an offline collection of state-action sequences, with associated rewards. In [20] it was shown that DDPMs exhibit good generalization and long-horizon planning abilities, even for sparse rewards. Furthermore, they showed that DDPMs can generate novel trajectories by locally stitching together sequences from the training distribution. Moreover, Ajay et al. [21] found that DDPMs can generate plans that meet multiple conditions even though these conditions were never met simultaneously in the training data. These properties of DDPMs make them potentially powerful and versatile planners. Real-world demonstrations of using DDPMs for robot control have shown promising results [22, 23, 24], which inspires this work to further investigate the novel aspects of using DDPMs to solve robotic planning problems. Carvalho et al. [24] propose a motion planner in the joint space using DDPMs. They use expert examples produced by an optimal motion planning algorithm to train a DDPM. During inference, they use the gradients of a cost function to bias the samples towards regions of low cost at every denoising step. Saha et al. [25] expand on this work by using an ensemble of cost functions to create a planner that can generate collision-free trajectories in a variety of environments. Chi et al. [22] use DDPMs to directly learn a policy that takes visual observations as its input and produces an action sequence as an output. In their work, they introduce closed-loop planning with DDPMs by implementing a receding-horizon control scheme. DDPM can also be used for safety-critical planning in combination with control barrier functions [26] to generate optimal and safe plans [27, 28].
@@
    A DDPM-based path planner for a robotic end-effector, not relying on a dynamic model of the robot and using only synthetic and low quality demonstrations;
@@
    Experimental validation on a 7-DoF robot manipulator, comparing five different conditional-sampling strategies.
@@ 3 Denoising Diffusion Planner
In this work, each path is discretized into a sequence of poses equally spaced in time, although the methodology applies to any parametrisations. The term “trajectory" is often used in literature for time-parameterized paths, and therefore we will use “path" and “trajectory" interchangeably from now on.
We indicate a path with $\bm{\tau}=\left(H_{0},H_{1},...,H_{T}\right)$, where $H$ is an element of the Lie-group SE(3) [32] which encodes the position and orientation of a frame with respect to a chosen reference frame. In particular, a pose $H_{t}$ can be described by a matrix that is composed of a rotation matrix $R_{t}$, encoding the orientation, and a vector $\bm{p}_{t}$, encoding the position [32]:
@@ 4 Experimental Design
Our code is entirely accessible at the following location: https://github.com/MSNikken/diffusion
@@ 4.2 Real-world Experiments
where
$d_{SE(3)}\left(H_{t},H_{g}\right)=\lVert\bm{p}_{t}-\bm{p}_{g}\rVert^{2}_{2}+\lVert LogMap\left(\bm{R}^{T}_{t}\bm{R}_{g}\right)\rVert^{2}_{2}.$
The $LogMap(\cdot)$ is a function from Lie Theory [32] that can be used to quantify the difference between two orientations. The collision cost is given by:
@@
with the weighting factors $w_{ee}$ and $w_{c}$. Thirdly, we experiment with a novel way of sampling through combining classifier-free sampling with cost-guided sampling. In total, this results in five combinations for conditional sampling:
  1. (i)
    Classifier-free guidance with dense rewards;
  2. (ii)
    Classifier-free guidance with sparse rewards;
  3. (iii)
    Cost guidance using (16);
  4. (iv)
    Classifier-free guidance with dense rewards and cost guidance (using only the end-effector cost (14));
  5. (v)
    Classifier-free guidance with sparse rewards and cost guidance (using only the end-effector cost (14)).
@@ 5 Results and Discussions
In this section, we show the results achieved with DDP. In particular, we study the impact of the choice of horizon and returns when conditionally-generating paths (Section 5.1), the effect of repeatedly inpainting the goal state (Section 5.2), and the performance of the DDP in real-world robotics tasks with different conditional-sampling approaches (Section 5.3).
@@ 5.1 Impact of Horizon and Returns
Since the training dataset contains paths in the entire workspace, even inside obstacles, the avoidance of obstacles is entirely achieved through conditional sampling. When using classifier-free guidance, these conditions are learned from the training data, which only consists out of straight lines. The extent to which the planner can generate paths that avoid an obstacle in between the start pose and goal pose depends on the DDPM’s generalization capabilities, as well as on the hyperparameters used in inference. In particular, we focus our studies on two important hyperparameters: (i) the planning horizon and (ii) the (expected) return on which the sample is conditioned. We consider a simple scenario in which the DDP is used in open-loop configuration to inpaint a path between a starting pose and a goal pose, while an obstacle is positioned in between. Figure 5 shows simulations of the paths for various combinations of horizon and return values. For the model trained on sparse rewards, only a return value of zero is considered, since this is the only value that represents a collision-free path. Therefore, we effectively only sweep over the horizon in the sparse-reward case. From Figure 5, it is evident that in all configurations, coherent and nontrivial paths are generated. These paths start at the intended starting pose and tend towards the goal pose, even for horizons that are much longer than seen in training (see Table 1). The paths that were generated using sparse rewards are close to the obstacles (Figure 5(b)), whereas for dense rewards, the distance to the obstacle is strongly influenced by the rewards (Figure 5(a)).
@@ 5.3 Real-world Results
We performed experiments in two different scenarios: the first with a single obstacle (for which Figure 7 shows some frames); and the second with two obstacles, which forces the robot to generate more complicated 3D paths to avoid collisions. Videos of our experiments can be found here. We compare five different conditional-sampling strategies as described in Section 4.2. The performed experiments confirmed that the DDP, trained only on low-quality synthetic data, is capable of repeatedly generate collision-free path in both test scenarios with different state and goal poses. The DDP demonstrated to be an extremely flexible framework that can exploit different configurations and conditional sampling strategies, as discussed in the previous section.
We list some observations that we consider significant and potentially useful for practitioners.
With reference to Figure 7, the DDP trained with sparse rewards (sparse) resulted in paths much closer to the obstacles compared to the DDP trained dense rewards (dense), regardless of the inclusion of cost guidance (cg). However, in the second scenario with two obstacles, using only classifier-free guidance with dense rewards often results in lingering in place before crossing from one side of the smaller obstacle to the other. It is worth mentioning that our reward functions, i.e., sparse and dense, only penalizes collisions or distance to obstacles, respectively, and reaching the goal state is controlled by the fixing the value of the final state(s) of the generated trajectory (see Section 5.2). Thus, except when utilizing cost guidance, the DDP is not explicitly conditioned to generate the shortest collision-free path from the initial to the goal state, but rather to generate an “acceptable", collision-free path.
We observed that the DDP with classifier-free guidance with dense rewards can be improved by using the cost function guidance (dense + cg). In particular, in the scenario with two obstacles, the DDP without cost function guidance would often linger in place to avoid moving closer to the obstacles. Adding the cost guidance allows for solving this shortcoming by pushing the poses to be close to the goal pose.
@@ 6 Conclusion and Future Work
In this work, we Denoising Diffusion Planner (DDP), an approach to design a path planner for a robotic end effector using DDPMs that does not use expert demonstrations, but only low-quality and synthetic data. Trained only with straight lines, the DDP can produce complex paths that avoid obstacles by utilizing the generalization capabilities of the DDPM and different conditional-sampling strategies. Additionally, deploying the DDP in a closed-loop receding-horizon control scheme and continuously update the plan allows for reaching difficult targets that cannot be reached in an open-loop configuration.